# Edit As You Wish: Video Caption Editing with Multi-grained User Control

### v2

@@ B.1. Automatic Dataset Construction
B.2. VATEX-EDIT Test set Quality
In the VATEX-EDIT dataset, it is worth noting that only the “del, pos, -, shorten description at specified positions” command uses the auto-constructed sentences as ground-truth captions in order to control the sentence structure. Meanwhile, the data samples of other commands utilize human-annotated sentences from the original VATEX dataset as ground-truth captions. To assess the quality of the auto-constructed sentences, we evaluate the test set of VATEX-EDIT through human and ChatGPT evaluations. For the ChatGPT555https://openai.com/blog/chatgpt evaluation, we design suitable prompts (depicted in Figure 12) to guide ChatGPT to judge whether a sentence is fluent or not. We randomly sample 100 ground-truth sentences on the test set and calculate the fluency rate obtained by ChatGPT. For human evaluation, we recruit 20 crowd workers to rate the fluency score (ranging from 1 to 5) of 200 randomly sampled ground-truth sentences. As Table 7 shows, the automatically constructed ground-truth sentences of the “del, pos, -” command are as fluent as the human-annotated ground-truth sentences of other commands (87% vs 91% and 4.40 vs 4.37), which indicates the high quality of the VATEX-EDIT test set.
@@ B.2. VATEX-EDIT Test set Quality
Appendix C EMMAD-EDIT Dataset Details
We manually collect the high-quality dataset EMMAD-EDIT in the E-commerce domain based on the Chinese E-commerce video captioning dataset E-MMAD. The data sample in E-MMAD dataset consists of a product video, an advertising video description, and additional information including video titles and structure attributes. It collects 120,984 videos with average duration of 30.4 seconds and the annotated Chinese description length is 67 words on average. The characteristics of long videos and long captions make it suitable for building a challenging VCE dataset. During annotation, we select data samples from E-MMAD dataset with relatively long descriptions. In addition to videos and descriptions, we also provide product structure information and video titles for reference.
C.1. EMMAD-EDIT dataset statistics
@@ D.1. Position Accuracy Design
The POS-Acc metric can align two variable-length text sequences in positional indexes, resulting in related spans {} in aligned to [MASK] tokens in . We count the percentage of correct samples that insert/remove new content in given position .
@@ E.2. From Triplets to Natural Language
The triplet command in our annotated datasets can be easily converted to natural languages to support more diverse scenarios and applications. For example, using natural language to explore the capability of LLMs in the VCE task (A.2). Specifically, we design a series of templates shown in Table 9 to convert the triplet command to natural language. We will also release the two benchmark datasets with user commands both in the triplet and natural language format to benefit the community.
@@
E.3. From Handwriting-revision Traces to Triplets
Appendix F Future Directions

### v3

@@ Appendix C EMMAD-EDIT Dataset Details
C.1. EMMAD-EDIT dataset statistics
@@ D.1. Position Accuracy Design
The pseudo-code is presented in Figure 17, which can align two variable-length text sequences in positional indexes, resulting in related spans {} in aligned to [MASK] tokens in . We count the percentage of correct samples that insert/remove new content in given position .
@@ E.2. From Triplets to Natural Language
The triplet command in our annotated datasets can be easily converted to natural languages to support more diverse scenarios and applications. For example, using natural language to explore the capability of LLMs in the VCE task (A). Specifically, we design a series of templates shown in Table 10 to convert the triplet command to natural language. We will also release the two benchmark datasets with user commands both in the triplet and natural language format to benefit the community.
E.3. From Handwriting-revision Traces to Triplets
Appendix F Future Directions
@@
Appendix G Social Impacts
The proposed Video Caption Editing (VCE) task addresses significant challenges in the realm of video content management, offering far-reaching social implications.
Enhanced Accessibility. By allowing for multi-grained user control over video captions, our system can significantly improve accessibility for individuals with disabilities. Users who rely on captions for understanding video content, such as the hearing-impaired, can benefit from more precise and customizable descriptions tailored to their specific needs.
E-commerce and Marketing. The integration of our VCE system into e-commerce platforms can revolutionize how product videos are presented. By enabling dynamic and detailed video descriptions that cater to individual consumer preferences, businesses can enhance user engagement and improve product understanding.
User Empowerment and Creativity. The ability to edit video captions interactively empowers users to tailor content to their specific and creative desires. This democratizes content creation, allowing individuals and small creators to produce high-quality, customized video content without requiring extensive technical expertise.